### v1

@@ 1. Introduction
Let ${\mathscr{M}}_{g,n}$ denote the moduli stack of smooth projective curves of genus $g$ and $n$-marked points, let ${\mathscr{A}}_{g}$ denote the moduli stack of principally polarized $g$-dimensional abelian varieties, and let ${\mathscr{A}}_{g,n}$ the $n$-fold fibre product of the universal family over ${\mathscr{A}}_{g}$. We work over an algebraically closed base field $k$. In this setting (and over $k={\mathbb{C}}$), we have the following results due to Harer-Zagier [15] and Harder [14]:
@@
(We take $\operatorname{sign}x$ to be $1$ if $x\geq 0$ and $-1$ otherwise.) ∎
In this article, we generalize this corollary by showing that the sign of the Euler characteristic of any smooth closed irreducible substack ${\mathscr{X}}$ of one the previous stacks satisfies the same condition. In fact, we can strengthen this as follows. Given a Deligne-Mumford stack defined over $k$, let $\operatorname{Perv}({\mathscr{X}})$ the category of perverse sheaves. In the following, this will mean one of two things. If $k={\mathbb{C}}$, one can consider the underlying analytic Deligne-Mumford stack and perverse sheaves with ${\mathbb{Q}}$ coefficients (in the analytic topology). For an arbitrary arbitrary field $k$ with $char(k)=p$, one can consider the category of $\ell$-adic perverse sheaves (for some $\ell\neq p$). In this setting and when $char(k)=0$, one can define the Euler-Satake characteristic of a perverse sheaf (as a rational number). The precise definition of $\operatorname{Perv}({\mathscr{X}})$ and the notion of Euler characteristic for an object of this category is given in section 3. Let us start by stating a special case of our main results.
@@
Let ${\mathscr{X}}$ be one of ${\mathscr{M}}_{g,n}$ (with $02g-2+n>$ 0) or ${\mathscr{A}}_{g,n}$, and suppose $char(k)=0$. Then for ${\mathcal{K}}\in\operatorname{Perv}({\mathscr{X}})$
@@
Take ${\mathcal{K}}=\iota_{*}{\mathbb{Q}}_{{\mathscr{X}}}[\dim{\mathscr{X}}]$ where $\iota$ is the inclusion of ${\mathscr{X}}$. This is a perverse sheaf by [18, chap III, lemma 6.5]. ∎
@@
Essentially by definition, ${\mathscr{X}}$ will embed as closed smooth substack of ${\mathscr{A}}_{g}$ for some $g$. ∎
In particular, this theorem gives a simple proof recovering the sign of the Euler characteristic of ${\mathscr{M}}_{g,n}$ and ${\mathscr{A}}_{g,n}$. We expect that the last corollary should hold for arbitrary Shimura varieties.
@@
By using the moduli spaces $\prescript{}{N}{{\mathscr{A}}_{g,n}}$ with level $N\geq 3$ structures, we reduce to an analogous statement about varieties rather than stacks. To check this version of the theorem, we verify that the log cotangent bundle of a product of toroidal compactifications
@@
is nef. This is reduced to a formula of Faltings and Chai, which expresses the log cotangent bundle of each factor as the second symmetric power of a Hodge bundle, and the Fujita-Kawamata semipositivity theorem for the latter. The above theorem now follows from the nefness result together with the log Dubson-Kashiwara theorem and an inequality of Fulton-Lazarsfeld. This last step of the proof is similar to, and inspired by, arguments used in recent work on the Hopf-Singer and related conjectures (see [4, 23]). We also note that the idea of using the Dubson-Kashiwara formula to show positivity of Euler charactersitic of perverse sheaves first appears (in the context of semi-abelian varieties) in the work of Franecki-Kapranov ([11]).
Our second result concerns moduli of curves. It is convenient to work with the moduli stack ${\mathscr{M}}_{g,n}^{c}\subset\overline{{\mathscr{M}}_{g,n}}$ of genus $g$ stable curves of compact type with $n$ marked points. Compact type means that the dual graph is a tree. In particular, ${\mathscr{M}}_{g,n}\subset{\mathscr{M}}_{g,n}^{c}$. Our second main theorem is the following:
@@
Let ${\mathscr{X}}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$ where $g_{i}>1$ or $g_{i}=1,n_{i}>0$, and suppose $char(k)=0$. If ${\mathcal{K}}$ is a perverse sheaf on ${\mathscr{X}}$, then
@@
which agrees with the Torelli map on ${}_{N}{\mathscr{M}}_{g}$. It easy to see that $\tau$ is a locally closed embedding of ${}_{N}{\mathscr{M}}_{g,n}$, but it is not an embedding of the whole space. So Theorem 1.6 does not follow immediately from Theorem 1.5. On the other hand, Theorem 1.5 does apply to perverse sheaves supported on the image $Y=\tau({}_{N}{\mathscr{M}}_{g,n}^{c})$. Now setting $X={}_{N}{\mathscr{M}}_{g,n}^{c}$ and $U={}_{N}{\mathscr{M}}_{g,n}$, we conclude the proof with the help of the following technical result proved in section 6 using Beilinson’s gluing construction (for perverse sheaves) and a deformation to the normal cone technique.
@@
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{j}}$$\scriptstyle{j}$${Y}$
where $j,\tilde{j}$ are open immersions, and $f$ is projective and birational, and $Z=X-U$ is a Cartier divisor. Suppose that for all $K^{\prime}\in\operatorname{Perv}(Y)$ and $K^{\prime\prime}\in\operatorname{Perv}(Z)$ we have $\chi(Y,K^{\prime})\geq 0$ and $\chi(Z,K^{\prime\prime})\geq 0$. Then for all $K\in\operatorname{Perv}(X)$, we have $\chi(X,K)\geq 0$.
In the final section of this paper, we show that the analog of Theorem 1.2 fails in positive characteristic. Let us now view $\prescript{}{N}{{\mathscr{A}}}_{g}$ as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$, and let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes\bar{{\mathbb{F}}}_{p}$ denote the fibre over $\bar{{\mathbb{F}}}_{p}$. We first observe that the ($\ell$-adic) Euler characteristic is indepenpent of $p\nmid N$ (see 9.5).
@@
This is closely related to the fact, first observed by Moret-Bailly [25], that the Fujita-Kawamata theorem fails in positive characteristic. See Remark 9.7. We note that we work with moduli with level structure, because it is not clear how to define the Euler-Satake characteristic of a Deligne-Mumford stack in characteristic $p$.
Finally, we briefly outline the contents of the various sections. The second section recall some background results which will be used in the following sections. In particular, we recall a positivity result of Fulton-Lazarsfeld, a log version of the Dubson-Kashiwara formula, and apply this to prove semi-positivity of certain Hodge bundles. We also recall a fundamental result of Beilinson on gluing perverse sheaves. In section 3, we give the definition of the Euler-Satake characteristic of construcitble sheaves on Deligne-Mumford stacks (following [5]), and recall some facts on the perverse sheaves on stacks. In particular, we observe that the constant sheaf on a local complete intersection stack is a perverse sheaf. In section 4, we prove the signed Euler characteristic property for ${}_{N}{\mathcal{A}}_{g}$ and their products. In section 5, we prove analogous assertions for the universal family over ${}_{N}{\mathcal{A}}_{g}$, and apply this to prove Theorem 1.5. In section 6, we prove Theorem 1.7. In section 7, we construct an extension of the Torelli map to ${\mathscr{M}}^{c}_{g,n}$. Finally, these results are applied in section 8 to prove Theorem 1.6 and finish the proof of Theorem 1.2. In section 9, we discuss the aforementioned results in characteristic p.
Acknowledgements: The first author thanks Laurentiu Maxim and Botong Wang from introducing him to some of these ideas in connection with the Hopf-Singer conjecture.
@@ 2. Background
In this section, we recollect some standard results which will be useful in the following.
@@ 2.1. Fulton-Lazarsfeld for nef bundles
Let $X$ be a smooth projective variety, and $E$ be a vector bundle. A conical subset of $E$ is a Zariski closed subset invariant under the natural ${\mathbb{G}}_{m}$-action. In this setting, one has the following positivity of intersection products result due to Fulton-Lazarsfeld.
@@
Suppose $E$ is nef of rank $e$. Then for any conical subset $C\subset E$ of dimension $e$,
@@
where $\langle C,X\rangle$ is the degree of the zero cycle given by the intersection of $C$ with the zero section.
@@ 2.2. Log Dubson-Kasiwara
Let $X$ be a smooth complex variety of dimension $d$, and ${\rm D}^{b}_{c}(X)$ denote the bounded derived category of constructible sheaves (of ${\mathbb{C}}$-vector spaces) on $X$. Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(X)$, there is a well-defined cycle $CC({\mathcal{K}})\in Z^{d}(T^{*}{X})$, where $Z^{d}(T^{*}(X)$ is the group of cycles of codimension $d$. This construction satisfies the following well known properties which we collect below:
@@
    If ${\mathcal{P}}$ is a perverse sheaf on $X$, the $CC({\mathcal{P}})$ is an effective cycle.
@@
    Here the right hand side is the degree of the zero cycle given by the intersection of the characteristic cycle and the zero section.
In ([29]), Wu and Zhou give an extension of the Dubson-Kashiwara formula to the setting of log cotangent bundles. Suppose now that $(X,D)$ is a pair where $X$ is a smooth projective variety and $D$ is a simple normal crossings divisor. In this case, one have the standard exact sequence:
$0\rightarrow\Omega^{1}_{X}\rightarrow\Omega^{1}_{X}(\log D)\rightarrow{\mathcal{O}}_{D}\rightarrow 0$
@@
Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(U)$, let $\overline{CC({\mathcal{K}})}$ denote its closure in $T^{*}X(\log D)$. This is a conical subset of dimension $d$. In [29, theorem 1.6], the authors show that
@@ 2.3. Kähler hyperbolicity
Given a Kähler manifold $X$, the universal cover $\pi:\tilde{X}\to X$ carries an induced Kähler metric. The manifold $X$ is Kähler hyperbolic if the pullback of the Kähler form $\pi^{*}\omega=d\alpha$, where $\alpha$ has bounded norm with respect to the induced metric. We have the following examples [13, ex 0.3A]:
@@ 2.4. Log structures
It is convenient to formulate some results in the language of log structures. We will not define them here, but instead refer the reader to Kato’s paper [16] for details. Let $k$ denote a field of characteristic 00. Given a smooth $k$-variety $Y$, a divisor $D\subset Y$ with reduced normal crossings induces a log structure [16, ex 1.5], which we denote by $(Y,D)$. We refer to a log structure of this type as having normal crossing type or just nc. If $f:X\to Y$ is a morphism of smooth varieties, with reduced divisors with normal crossings $D\subset Y$ and $E\subset X$ such that $f^{-1}D\subseteq E$, then we get morphism of nc log schemes $f:(X,E)\to(Y,D)$. We say $f$ is log smooth if it is a smooth morphism of log schemes [16, §3.3]. This implies that

### v2

@@ 1. Introduction
Let ${\mathscr{M}}_{g,n}$ denote the moduli stack of smooth projective curves of genus $g$ and $n$-marked points, let ${\mathscr{A}}_{g}$ denote the moduli stack of principally polarized $g$-dimensional abelian varieties, and let ${\mathscr{A}}_{g,n}$ the $n$-fold fibre product of the universal family over ${\mathscr{A}}_{g}$. We work over an algebraically closed base field $k$. In this setting, when $k={\mathbb{C}}$, we have the following results due to Harer-Zagier [hz] and Harder [harder]:
@@
(We take $\operatorname{sign}x$ to be $1$ if $x\geq 0$ and 1-1 otherwise.) ∎
In this article, we generalize this corollary by showing that the sign of the Euler characteristic of any smooth closed irreducible substack $\mathscr{X}$} of one the previous stacks satisfies the same condition. In fact, we can strengthen this as follows. Given a Deligne-Mumford stack defined over $k$, let $\operatorname{Perv}({\mathscr{X}})$ be the category of perverse sheaves. In the following, this will mean one of two things. If $k={\mathbb{C}}$, one can consider the underlying analytic Deligne-Mumford stack and perverse sheaves with $\mathbb{Q}$} coefficients (in the analytic topology). For an arbitrary field $k$ with $char(k)=p$, one can consider the category of $\ell$-adic perverse sheaves (for some $\ell\neq p$). In this setting and when $char(k)=0$, one can define the Euler-Satake characteristic of a perverse sheaf (as a rational number). The precise definition of $\operatorname{Perv}({\mathscr{X}})$ and the notion of Euler characteristic for an object of this category is given in section 2. Let us start by stating one of our main results.
@@
Let $\mathscr{X}$} be one of ${\mathscr{M}}_{g,n}$ (with $2g-2+n\geq 3$) or ${\mathscr{A}}_{g,n}$, and suppose $char(k)=0$. Then for ${\mathcal{K}}\in\operatorname{Perv}({\mathscr{X}})$
@@
Another motivation for this theorem comes from the work of Liu, Maxim and Wang [LMW], the author and Wang [aw], and Deng and Wang [dw]. Similar inequalities were obtained in [aw] and [dw] (and conjectured in [LMW]) for perverse sheaves on a projective manifold that admits a nonpositively curved Kähler metric or a large local system. A local system is large if it restricts nontrivially to all curves mapping nontrivially to the variety. These conditions certainly hold for the orbifold ${\mathscr{A}}_{g}$ since it is a locally symmetric space, and for ${\mathscr{M}}_{g}$ since it essentially embeds into it. So theorem 1.2 should be viewed as an extension of the aforementioned results to these sorts of noncompact spaces.
@@
Take ${\mathcal{K}}=\iota_{*}{\mathbb{Q}}_{{\mathscr{X}}}[\dim{\mathscr{X}}]$ where $\iota$ is the inclusion of $\mathscr{X}$}. This is a perverse sheaf by [kw, chap III, lemma 6.5]. ∎
@@
Essentially, by definition, $\mathscr{X}$} will embed as a closed smooth substack of ${\mathscr{A}}_{g}$ for some $g$. ∎
In particular, this theorem gives a simple proof that recovers the sign of the Euler characteristic of ${\mathscr{M}}_{g,n}$ and ${\mathscr{A}}_{g,n}$. We expect that the last corollary should hold for arbitrary Shimura varieties. For compact Shimura varieties, this follows from [aw].
@@
Using moduli spaces $\prescript{}{N}{{\mathscr{A}}_{g,n}}$ with level structures $N\geq 3$, we reduce to an analogous statement about varieties rather than stacks. To check this version of the theorem, we verify that the log cotangent bundle of a product of toroidal compactifications
@@
is nef. This is reduced to a formula of Faltings and Chai, which expresses the log cotangent bundle of each factor as the second symmetric power of a Hodge bundle, and the Fujita-Kawamata semipositivity theorem for the latter. The above theorem now follows from the nefness result together with the log Dubson-Kashiwara theorem and an inequality of Fulton-Lazarsfeld. This last step of the proof is similar to and inspired by the arguments used in the papers [aw, dw, LMW] mentioned above. We also note that the idea of using the Dubson-Kashiwara formula to show positivity of the Euler characteristic of perverse sheaves first appears (in the context of semi-abelian varieties) in the work of Franecki-Kapranov ([fk]).
Our second result concerns moduli of curves. It is convenient to work with the moduli stack ${\mathscr{M}}_{g,n}^{c}\subset\overline{{\mathscr{M}}_{g,n}}$ of genus $g$ stable curves of compact type with $n$ marked points, and its substack ${\mathcal{M}}_{g,n}^{sc}$. [^1] By definition, this is the open substack of compact type curves such that any genus zero vertex is only connected to higher genus vertices. See Section 6. One has ${\mathscr{M}}_{g,n}\subset{\mathscr{M}}_{g,n}^{sc}\subset{\mathscr{M}}_{g,n}^{c}$. Our second main theorem is the following:
@@
Let ${\mathscr{X}}={\mathscr{M}}_{g_{1},n_{1}}^{sc}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{sc}$ where $2g_{i}-2+n\geq 3$, and suppose $char(k)=0$. If $\mathcal{K}$} is a perverse sheaf on $\mathscr{X}$}, then
@@
which agrees with the Torelli map on $_{N}{\mathscr{M}}_{g,n}$. An application of Theorem 1.5 shows that the conclusion of the previous theorem does hold for $Y=\tau({}_{N}{\mathscr{M}}_{g,n}^{sc})$. On the other hand, since $\tau$ restricted to $_{N}{\mathscr{M}}_{g,n}$ is not an affine open embedding, Theorem 1.6 does not immediately follow. Our strategy is to induct on the boundary strata for which the stack ${\mathcal{M}}_{g,n}^{sc}$ is a useful technical tool. The following technical result, proved in section 5 using Beilinson’s gluing construction (for perverse sheaves) and a deformation to the normal cone technique, allows us to deduce Theorem 1.6 via induction on boundary strata and Theorem 1.5.
@@
[…]
where $j,\tilde{j}$ are open immersions, $f$ is birational, and Z=XUZ=X-U is a Cartier divisor. Suppose that for all $K^{\prime}\in\operatorname{Perv}(Y)$ and $K^{\prime\prime}\in\operatorname{Perv}(Z)$ we have $\chi(Y,K^{\prime})\geq 0$ and $\chi(Z,K^{\prime\prime})\geq 0$. Then for all $K\in\operatorname{Perv}(X)$, we have $\chi(X,K)\geq 0$.
For all but the final section of the paper, the ground field $k$ will have characteristic 0. In the final section, we show that the analog of Theorem 1.2 fails in positive characteristic. Let us now view $\prescript{}{N}{{\mathscr{A}}}_{g}$ as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$, and let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes\bar{{\mathbb{F}}}_{p}$ denote the fibre over $\bar{{\mathbb{F}}}_{p}$. We first observe that the ($\ell$-adic) Euler characteristic is independent of $p\nmid N$ (see 8.5).
@@
This is closely related to the fact, first observed by Moret-Bailly [mb], that the Fujita-Kawamata theorem fails in positive characteristic. See Remark 8.7. We note that we work with moduli with level structure, because it is not clear how to define the Euler-Satake characteristic of a Deligne-Mumford stack in characteristic $p$.
Finally, we briefly outline the contents of the various sections. The second section recalls some background results which will be used in the following sections. In particular, we recall a positivity result of Fulton-Lazarsfeld, a log version of the Dubson-Kashiwara formula, and apply this to prove semi-positivity of certain Hodge bundles. We also give the definition of the Euler-Satake characteristic of constructible sheaves on Deligne-Mumford stacks (following [Beh]), and recall some facts on the perverse sheaves on stacks. In particular, we observe that the constant sheaf on a local complete intersection stack is a perverse sheaf. In section 3, we prove the signed Euler characteristic property for $_{N}{\mathcal{A}}_{g}$ and their products. In section 4, we prove analogous assertions for the universal family over $_{N}{\mathcal{A}}_{g}$, and apply this to prove Theorem 1.5. In section 5, we prove Theorem 1.7. In section 6, we construct an extension of the Torelli map to ${\mathscr{M}}^{c}_{g,n}$. Finally, these results are applied in section 7 to prove Theorem 1.6 and finish the proof of Theorem 1.2. In section 8, we discuss the aforementioned results in characteristic p.
Acknowledgements: The first author thanks Laurentiu Maxim and Botong Wang for introducing him to some of these ideas in connection with the Hopf-Singer conjecture.
@@ 2. Background
In this section, we recall some standard results which will be useful in the following.
@@ 2.1. Fulton-Lazarsfeld for nef bundles
Let $X$ be a smooth projective variety, and $E$ be a vector bundle of rank $e$. Since the zero section $\iota:X\hookrightarrow E$ is a regular embedding, we have a Gysin map $\iota^{!}:CH_{e}(E)\to CH_{0}(X)$. Let
$\langle C,X\rangle=\operatorname{deg}\iota^{!}[C]$
for an $e$-cycle $C$ on $E$. This should be viewed as the intersection number of $C$ with the zero section. A conical subset $C\subset E$ is a Zariski closed subset invariant under the natural ${\mathbb{G}}_{m}$-action. In this setting, one has the following positivity of intersection products result due to Fulton-Lazarsfeld.
@@
Suppose that $E$ is a nef vector bundle of rank $e$. Then for any conical subset $C\subset E$ of dimension $e$,
@@ 2.2. Log Dubson-Kasiwara
Suppose that $k={\mathbb{C}}$. Let $X$ be a smooth variety of dimension $d$, and ${\rm D}^{b}_{c}(X)$ denote the bounded derived category of constructible sheaves (of $\mathbb{C}$}-vector spaces) on $X$. Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(X)$, there is a well-defined cycle $CC({\mathcal{K}})\in Z^{d}(T^{*}{X})$, where $Z^{d}(T^{*}(X))$ is the group of cycles of codimension $d$. This construction satisfies the following well known properties which we collect below:
@@
    If $\mathcal{P}$} is a perverse sheaf on $X$, then $CC({\mathcal{P}})$ is an effective cycle.
@@
In ([wz]), Wu and Zhou give an extension of the Dubson-Kashiwara formula to the setting of log cotangent bundles. Suppose now that $(X,D)$ is a pair where $X$ is a smooth projective variety and $D$ is a simple normal crossings divisor. In this case, one have the standard exact sequence:
$0\rightarrow\Omega^{1}_{X}\rightarrow\Omega^{1}_{X}(\log D)\rightarrow{\mathcal{O}}_{D}\rightarrow 0.$
@@
Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(U)$, let $\overline{CC({\mathcal{K}})}$ denote its closure in $T^{*}X(\log D)$. This is a conical subset of dimension $d$. In [wz, theorem 1.6], the authors show that
@@
Putting these results together yields the following fact needed below.
Lemma 2.2.
Proof.
$\chi(A,P)=\langle\overline{CC(P)},A\rangle$
Therefore it is enough to show that this intersection number is nonnegative. Since the log cotangent bundle is nef. Theorem 2.1 yields the desired nonnegativity.
We will need the following as well.
Lemma 2.3.
Let $X_{1}$ and $X_{2}$ be smooth projective varieties with snc divisors $E_{i}$, such that $\Omega_{X_{i}}^{1}(\log E_{i})$ are nef. Then $\Omega_{X_{1}\times X_{2}}^{1}(\log p_{1}^{*}E_{1}+p_{2}^{*}E_{2})$ is nef, where $p_{i}:X_{1}\times X_{2}\to X_{i}$ are the natural projections.
Proof.
This follows from [lazarsfeld, theorem 6.2.12] and the fact that
@@ 2.3. Kähler hyperbolicity
Given a Kähler manifold $X$, the universal cover $\pi:\tilde{X}\to X$ carries an induced Kähler metric. The manifold $X$ is Kähler hyperbolic if the pullback of the Kähler form $\pi^{*}\omega=d\alpha$, where $\alpha$ has bounded norm with respect to the induced metric. We have the following examples [gromov, ex 0.3A]:
@@ 2.4. Log structures
It is convenient to formulate some results in the language of log structures. We will not define them here, but instead refer the reader to Kato’s paper [kato] for details. Given a smooth $k$-variety $Y$, a divisor $D\subset Y$ with reduced normal crossings induces a log structure [kato, ex 1.5], which we denote by $(Y,D)$. We refer to a log structure of this type as having normal crossing type or just nc. If $f:X\to Y$ is a morphism of smooth varieties, with reduced divisors with normal crossings $D\subset Y$ and $E\subset X$ such that $f^{-1}D\subseteq E$, then we get a morphism of nc log schemes $f:(X,E)\to(Y,D)$. We say $f$ is log smooth if it is a smooth morphism of log schemes [kato, §3.3]. This implies that